OpenStudy (cookie_monster):

HELP PLEASE!!!! x + y = 7- What is the y and x intercept for this?

OpenStudy (solomonzelman):

\[x + y = 7~~~~~~~~~subtract~~x~~from~~both~~sides.~~~~->~~~~y=-x+7\]\[compare~~\to~~~y=mx+b,~~~~where~~m~~is~~slope,~~and~~b~~is~~y-intercept.\]\[y=(-1)x+(-7),~~~~m=-1~~~~b=-7\]

OpenStudy (cookie_monster):

This is a little confusing...@SolomonZelman

OpenStudy (anonymous):

He just means you want to get it into the form y=mx+b. You want to get your y on a side by itself. c: So you subtract the x from both sides, and you're left with y=-x+7. Since this form is called slope-intercept, we know that, or we learn that m is our slope, and b is our y-intercept. Then we can graph it, hang on...

OpenStudy (anonymous):